OpenStudy (anonymous):

Determine the values of k which will make the roots of the equation x^2-kx-k=0 equal.

OpenStudy (amistre64):

have you tried applying the quadratic formula?

OpenStudy (amistre64):

if the roots are equal then the sqrt needs to be zero ....

OpenStudy (anonymous):

There are to answer, right? Is it 0, -4?

OpenStudy (amistre64):

that does look correct to me k^2 +4k = 0 k(k +4) = 0 yep

OpenStudy (anonymous):

So, the answer are 0 and -4?

OpenStudy (amistre64):

so long as k = 0 or -4 makes: k(k +4) = 0 a true statement, then yes